OpenStudy (anonymous):

Determine whether the given quadratic function has a minimum value or maximum value. Then find the coordinates of the minimum or maximum point. f(x) = x2 + 2x - 9

OpenStudy (anonymous):

Can someone help me?

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

Can someone help me???

OpenStudy (anonymous):

There are multiple ways of doing this. I will show you the method of making a perfect square. f(x) = x^2 + 2x - 9 f(x)=(x+1)^2 -10 as (x+1)^2=x^2+2x+1 (x+1)^2 is always positive or 0, at x=-1 it will be 0, at any other point it will take higher values. Thus the function will have its minimal value at x=-1, where it will take (-1+1)^2 -10=0-10=-10 So the x,y coordinates are (-1,-10). Minimal point

OpenStudy (anonymous):

@Andras I got Vertex: (−1,−10)

OpenStudy (anonymous):

Can anyone help me????

OpenStudy (anonymous):

Help????????

OpenStudy (anonymous):

What is wrong. U got the same thing I wrote. Make a graph to help you understand it.